# Task Prompt Vectors: Effective Initialization through Multi-Task Soft-Prompt Transfer

### v1

@@ 4.3 Combination of Task Prompt Vectors via Arithmetic Addition for Multi-Task Transfer
To answer this research question, we investigate the method of combination via addition on 15 task pair combinations from the set of NLU datasets (MNLI, QNLI, DBPedia, TREC Coarse, SST2, Yelp Polarity). We also evaluate combinations of task prompt vectors in a simulated limited data environment by providing 0 to 100 training examples before evaluation on the test set.
@@
Table 2: Test results of training T5-base model with random, single- and multi-task soft-prompt transfer (SPoT), multi-task ATTEMPT, and our task prompt vectors on 0-shot and 100-shots of data. We show the initialization with different combinations for natural language inference classification, topic classification, and sentiment classification. The subscript represents the standard deviation from the average. The best results are bold, while the second-best results are underlined. The * in the superscript represents that the results are statistically significant from the others, by two-sample Student’s t-test (Student, 1908) or Welch’s t-test (Welch, 1947).
@@
To evaluate how the combinations of task prompt vectors maintain their single-task performance, we conduct experiments of creating pair combinations from all of the source tasks (according to equation 4). After we create the task prompt vector combinations, we evaluate their performance on the individual source tasks that formed the task combination and find the best rescaling factor $\lambda$ via held-out validation sets. We aggregate the best-performing combinations in Figure 4. The full results from the experiment can be found in Appendix C in Figure 10. We can see from the results that most of the binary classification tasks retain their single-task performance on both of the tasks, which implies that task prompt vectors can be used for solving multi-task problems. In some cases, the single-task performance was kept only for a single source task, which leads us to the conclusion that certain combinations of task prompt vectors may be more suitable than others.
Now that we know how the combinations of task prompt vectors affect the performance across source tasks, we evaluate the combinations of task prompt vectors on a set of out-of-distribution target tasks from the same problem area. We chose two target tasks for inference classification (SciTail, SNLI), topic classification (AG News, Yahoo Answers), and sentiment classification (SST5, IMDB) and we keep the same set of source tasks. Results for SciTail, AG News, and IMDB are in Table 2; the full table with extended experiments is in Appendix D in Table 4.
@@
Table 3: Comparison of multi-task properties for compared methods. Task Prompt Vectors maintain high task modularity, multi-task performance, and are independent of the quality or number of pre-trained source soft-prompts.
Task prompt vector combinations can be used as an initialization for zero-shot and few-shot learning.
We conduct 0-shot and 100-shot evaluations of different initialization of prompt tuning on the set of target tasks. We compare initialization with randomly initialized soft-prompt, soft-prompt trained on single and multiple source tasks (this is an equivalent of soft-prompt transfer presented in Vu et al. (2022)), the multi-task ATTEMPT Asai et al. (2022) method, and a combination of task prompt vectors of both of the source tasks. From the results in Table 2 we can see that our combination of task prompt vectors can outperform the initialization with a single-task source soft-prompt on SciTail and IMDB datasets and the multi-task source soft-prompt only in the case of SciTail task. In some cases, the combination of task prompt vectors matched the SPoT baseline like in the case of AG News. This may be because the combination of DBPedia and TREC does not retain much information about TREC which could benefit the overall result.
We can also see that the ATTEMPT method is significantly underperforming when using a smaller set of pre-trained source soft-prompts. Another observation is that ATTEMPT performs better on the AG News task. This may be caused by using the original implementation of ATTEMPT, where authors instead of using textual labels (i.e., "entailment", "not entailment") used textual numbers as labels (i.e., "0", "1"), which made the model to predict numbers instead of specific words (unlike in the other methods).
While matching the results of full multi-task soft-prompt transfer (SPoT) training initialization of prompt tuning using Task Prompt Vectors combinations also retains high task modularity, which means that we can add new tasks without the necessity of training. Only in the case of the IMDB task, the SPoT baseline fine-tuned on both datasets performs better, however, it requires the training process to be fully repeated for each new task, resulting in higher computational costs. Table 3 compares attributes beneficial for multi-task training for SPoT, ATTEMPT, and Task Prompt Vectors methods. We can see that the SPoT method has low multi-task modularity because we need to re-train the source soft-prompt every time we change the set of source tasks. ATTEMPT, while having sufficient task modularity, depends heavily on the quality and number of source soft-prompts. Task Prompt Vectors have both of these attributes and also retain sufficient multi-task performance.
Figure 5: Test results of training T5-base model with random, single and multi-task soft-prompt transfer (SPoT), multi-task ATTEMPT, and our task prompt vectors combination on increasing numbers of shots of data. We can see that for SciTail and IMDB tasks combination of task prompt vectors outperforms single task transfer.
@@ 4.4 Additional Results: Few-Shot Comparison
In this section, we study how the increasing number of data affects the performance of prompt tuning on a target task initialized by a combination of task prompt vectors of similar source tasks. We keep the same experiment setup as in the previous section and further evaluate the soft-prompt initialization on 5,10,25,100,250,500 shots. We evaluate the topic classification tasks also for 500 shots since we are starting from 10 shots due to our sampling method. The results from this set of experiments can be seen in the few-shot plots in Figure 5.
From the results, we can see that the performance of the combination of task prompt vectors for SciTail and IMDB target tasks outperforms using a single-task initialization for multiple shots. We can also see that our method also outperforms the multi-task initialization for the SciTail dataset across all shots of data. The most significant difference in performance is in 10 shots for SciTail and in 5 shots for IMDB. This also means that the choice of the source tasks for transfer learning plays an important role in the initialization with task prompt vector combinations. Comparing the results from Figure 4 and Figure 5, if we choose a combination of tasks that maintains a significant amount of the source task performance (MNLI + QNLI and SST2 + Yelp), the few-shot performance of the task prompt vector combination tend to be higher than single-task transfer. The full results across more shots and more target tasks can be found in Appendix D in Figure 11.
@@ 5 Discussion
In Section 4.2, we study the properties of task prompt vectors. We first show that the task prompts and their corresponding task prompt vectors are close to orthogonal, by comparing their cosine similarities across multiple initializations in Figures 2 and 3. This may mean that the task prompts are in different places in the embedding space of the model and also that the task prompt vectors have different directions. Despite their near-orthogonality, task prompt vectors created from one initialization and applied to a different one maintain their performance for the majority of the observed tasks, as we can see from Table 1. The implication of this finding may mean that it is possible to combine are reuse different task prompt vectors from different initializations, therefore we can re-use pre-trained task prompt vectors for different tasks and use them in downstream scenarios like for initialization for further fine-tuning.
In Section 4.3, we focus mainly on the combination of task prompts via arithmetic addition. We show that combinations of certain task prompts maintain their source single-task performance (in Figure 4) and that the combinations of task prompt vectors can be used for initialization of prompt tuning (in Table 2) in simulated low resource setting on the set of target tasks. The combinations that retain most of their single-task performance on both source tasks also tend to have higher performance on 0 and 100 shots compared to the single-task soft-prompt transfer. The implication of findings from this set of experiments may be that we can use different combinations of task prompt vectors to gain even zero-shot multi-task behavior if we correctly choose the source tasks. We can combine multiple task prompt vectors and maintain multi-task performance on the source tasks, but the right task combinations need to be found (e.g., by evaluating on held-out validation sets).
In our last section of experiments (4.4) we extend our few-shot experiments to more samples of data, showing the development of performance across different initializations and different amounts of data. From the results in Figure 5, we can see that for the SciTail and IMDB datasets our task prompt vector initialization maintains its higher performance compared to the single-task soft-prompt transfer even with higher number of samples.
@@ 6 Conclusion
In our work, we introduce and investigate task prompt vectors as a method of multi-task transfer from prompt tuning. We show that the task prompt vectors are not dependent on random initialization and that the performance across different random initializations does not change significantly in the majority of observed source tasks. Additionally, we show that in some tasks the combination via arithmetic addition maintains the single-task performance. Finally, we show that certain combinations of task prompt vectors can be a better option for initialization in a simulated limited data environment for certain tasks while maintaining higher multi-task modularity than other methods.
In the future, we would like to extend our work by evaluating the cross-model performance of task prompt vectors. Moreover, task prompt arithmetic has the highest potential for improving the unlearning in PLMs by subtracting (negating) the task prompt vectors for the tasks that we want to unlearn. Such an option is enabled by introducing task prompt vectors, which it would not be possible with the existing state-of-the-art methods.
@@ Acknowledgements
The experiments in our work utilize only datasets in the English language with various characteristics (problem type, dataset size, number of classes), mostly because of the large amount of publicly available data covering a variety of NLU problems. At the same time, to direct our focus primarily on the evaluation of task prompt vectors, we utilize only monolingual models in the scope of our work.
Even though there are many other PLMs capable of conditional generation that beat T5 models in performance on various benchmarks, we focus our experiments on the T5-base model as it is commonly used as a representative model in many PEFT methods.
Finally, we focus on 3 common NLU problems (natural language inference, topic classification, and sentiment classification) that are commonly incorporated in NLU benchmarks and do not consider other NLU problems (e.g., question answering, slot tagging, acceptability classification). However, we find that our set of 3 common NLU problems each covering 4 different tasks, is enough to evaluate the properties of task prompt vectors.
@@
The experiments in this paper were conducted with publicly available datasets MNLI, QNLI, SciTail, SNLI, DBPedia, TREC Coarse, AG News, Yahoo Answers, SST2, Yelp Polarity, SST5, and IMDB, citing the original authors. MNLI, QNLI, and SST2 are part of the GLUE benchmark. As we were not able to determine the license for all used datasets, we have opted to use them as in a limited form as possible, adhering to the terms of use of the GLUE benchmark for all of the mentioned datasets. As the datasets are commonly used in other related works, and were published in scientific works that went through an established review process, we do not check for the presence of any offensive content as it was already removed by the authors of these publicly available datasets. In addition, we do not utilize any personally identifiable information or offensive content and we do not perform crowdsourcing in any form for data annotation. To our knowledge, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the field of Machine Learning (i.e., use of computational resources that are consuming energy and producing heat with indirect CO2 emission production). We follow the license terms for the T5-base model we use – all models and datasets allow their use as part of the research. As we perform conditional generation transform into the classification problem (generating only labels), we minimize the problem of generating offensive or biased content.
@@
The experiments in this paper require a significant amount of GPU computing resources as we train and evaluate 1 model over multiple random initializations (10) for different methods (4) and datasets (12). Overall the experiments including evaluations (which did not require training, but still used GPU resources for inference) and preliminary experiments (which are not reported in the scope of our work) were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh. Approximately, 1000 hours of computation performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 95.09 kgCO2eq of which 0 percent were directly offset. These estimations were conducted using the CodeCarbon Courty et al. (2024) python module. Whenever possible we tried to reduce the computational costs. Because our method is built upon the prompt tuning PEFT method, we always trained only a small part of the model parameters (76800 parameters, which is around 0.2% of the T5-base model parameters), and training the model fully will probably require more GPU hours and create more CO2 emissions.
@@ Appendix A Experimental setup: Further Details
We provide all of our configurations in the config directory of our repository. We set different hyperparameters for prompt tuning and the hyperparameters for zero- or few-shot evaluation do not differ much from the hyperparameters for prompt tuning. In general, we chose the maximum token length for labels by searching the dataset for the maximum token length (in our configs, we set default max_target_lenght to 128 if the dataset requires to generate sentences), for the inputs we pad the token sequences to 256 tokens with the max_target_lenght parameter. We use a learning rate of 0.3 for the AdamW optimizer, with weight decay of $1\times 10^{-5}$ and 500 warmup steps for 10 epochs (with an exception for the TREC Coarse dataset) with the batch size of 32. We evaluate, log and save after each 100 training steps and keep only the best model at the end of the training. In our configs, we set a number of tokens to 50, but in reality, Hugging Face peft library doubles the number for encoder-decoder models like T5.
@@ Appendix B Additional results: Task Prompt Vectors and Task Prompt Cosine Similarities
Figure 6: Comparisons of cosine similarities of task prompts fine-tuned on different tasks. Each heatmap represents a different task combination. We calculate the cosine similarities for all combinations of 10 random initializations omitting the combinations of random initializations where cosine similarity is equal to 1 (single-task comparisons). Each heatmap is represented as a single field in Figure 2 by averaging all values. The x and y axes represent the number of random initializations.
@@
Figure 8: Comparisons of average cosine similarities of task prompt vectors. The averages are calculated similarly to Figure 6 but with task prompt vectors created from different task prompts. Each heatmap represents a different task combination. We calculate the cosine similarities for all combinations of 10 random initializations omitting the combinations of random initializations where cosine similarity is equal to 1 (single-task comparisons). Each heatmap is represented as a single field in Figure 3 by averaging all values. The x and y axes represent the number of random initializations.

### v2

@@ 4.3 Combination of Task Prompt Vectors via Addition for Multi-Task Transfer
To answer this research question, we investigate the method of combination via addition on 15 task pair combinations from the set of NLU datasets (MNLI, QNLI, DBPedia, TREC, SST2, Yelp Polarity). We also evaluate combinations of task prompt vectors in a simulated limited data environment by providing 0 to 100 training examples before evaluation on the test set.
@@
Table 2: Test results of training T5-base model with random, single- and multi-task soft-prompt transfer (SPoT), multi-task ATTEMPT, and our task prompt vectors on 0-shot and 100-shots of data. We show the initialization with different combinations for natural language inference classification, topic classification, and sentiment classification. The subscript represents the standard deviation from the average. The best results are bold, while the second-best results are underlined. The * denotes statistical significance.
@@
To evaluate how the combinations of task prompt vectors maintain their single-task performance, we conduct experiments of creating pair combinations from all of the source tasks (according to equation 3). We aggregate the best-performing combinations in Figure 4. The full results from the experiment can be found in Appendix C in Figure 10. We can see from the results that most of the binary classification tasks retain their single-task performance on both of the tasks, which implies that task prompt vectors can be used for solving multi-task problems. In some cases, the single-task performance was kept only for a single source task, which leads us to the conclusion that certain combinations of task prompt vectors may be more suitable than others.
Additionally, we chose two target tasks for inference classification (SciTail, SNLI), topic classification (AG News, Yahoo Answers), and sentiment classification (SST5, IMDB) and we keep the same set of source tasks. Results for SciTail, AG News and IMDB are in Table 2; the full table with extended experiments is in Appendix D in Table 4.
@@
Table 3: Comparison of multi-task properties for compared methods. Task prompt vectors maintain high task modularity and multi-task performance, and are independent of the quality or number of pre-trained source soft-prompts.
Task prompt vector combinations can initialize zero-shot and few-shot learning.
We compare initialization with randomly initialized soft-prompt, soft-prompt trained on single and multiple source tasks (this is an equivalent of soft-prompt transfer presented in Vu et al. (2022)), the multi-task ATTEMPT Asai et al. (2022) method, and a combination of task prompt vectors of both of the source tasks. From the 0-shot and 100-shot results (Table 2), we can see that our combination of task prompt vectors can outperform the initialization with a single-task source soft-prompt on SciTail and IMDB datasets and the multi-task source soft-prompt only in the case of SciTail task. In some cases, the combination of task prompt vectors matched the SPoT baseline like in the case of AG News. This may be because the combination of DBPedia and TREC does not retain much information about TREC, which could benefit the overall result.
Figure 5: Test results of training T5-base model with random, single, and multi-task soft-prompt transfer (SPoT), multi-task ATTEMPT, and our task prompt vectors combination on increasing numbers of shots of data. We can see that for SciTail and IMDB tasks, a combination of task prompt vectors outperforms single task transfer.
We can also see that the ATTEMPT method is significantly underperforming when using a smaller set of pre-trained source soft-prompts. Another observation is that ATTEMPT performs better on the AG News task. This may be caused by using the original implementation of ATTEMPT, where authors, instead of using textual labels (i.e., "entailment", "not entailment"), used textual numbers as labels (i.e., "0", "1"), which made the model predict numbers instead of specific words.
While matching the results of full multi-task soft-prompt transfer (SPoT) training initialization of prompt tuning using task prompt vector combinations also retains high task modularity, which means that we can add new tasks without the necessity of training. Only in the case of the IMDB task does the SPoT baseline fine-tuned on both datasets perform better. However, it requires the training process to be fully repeated for each new task, resulting in higher computational costs. Table 3 compares attributes beneficial for multi-task training for SPoT, ATTEMPT, and task prompt vector methods. We can see that the SPoT method has low multi-task modularity because we need to re-train the source soft-prompt every time we change the set of source tasks. ATTEMPT, while having sufficient task modularity, depends heavily on the quality and number of source soft-prompts. Task prompt vectors have both of these attributes and also retain sufficient multi-task performance.
@@ 4.4 Additional Results: Few-Shot Comparison
In this section, we study how increasing the number of demonstration data affects the performance of prompt tuning on a target task initialized by a combination of task prompt vectors of similar source tasks. We keep the same experiment setup as in the previous section and further evaluate the soft-prompt initialization on 5, 10, 25, 100, 250, 500 shots. We also assessed the topic classification tasks for 500 shots since we started with 10 shots due to our sampling method.
From the results in Figure 5, we can see that the performance of the combination of task prompt vectors for SciTail and IMDB target tasks outperforms using a single-task initialization for multiple shots. We can also see that our method outperforms the multi-task initialization for the SciTail dataset across all shots of data.
Comparing the results from Figure 4 and Figure 5, if we choose a combination of tasks that maintains a significant amount of the source task performance (MNLI + QNLI and SST2 + Yelp), the few-shot performance of the task prompt vector combination tend to be higher than single-task transfer. The full results across more shots and more target tasks can be found in Appendix D in Figure 13.
@@ 5 Discussion
In Section 4.2, we showed that task prompts and their corresponding task prompt vectors are close to orthogonal, by comparing their cosine similarities across multiple initializations in Figures 2 and 3. Despite their near-orthogonality, task prompt vectors created from one initialization and applied to a different one maintain their performance for the majority of the observed tasks.
In Section 4.3, we showed that combinations of certain task prompts maintain their source single-task performance (in Figure 4) and that the combinations of task prompt vectors can be used for initialization of prompt tuning (in Table 2) in simulated low resource setting on the set of target tasks. From the results in Figure 5, we can see that, for the SciTail and IMDB datasets, our task prompt vector initialization maintains its higher performance compared to the single-task soft-prompt transfer even with the higher number of samples.
Theoretical implications and analysis.
It lies beyond the scope of our work to further deliver theoretical proofs for diverse properties of task prompt vectors, which we will leave for future work. However, we still discuss the implications that arise from the contributions of task prompt vectors.
Based on the observations from Section 4.2, we can re-use pre-trained task prompt vectors for different tasks and use them in downstream scenarios. Since task prompt vectors are independent of their initialization, we can also re-use pre-trained task prompt vectors shared on the internet (e.g., on a designated vector hub). Another implication that we can derive from these findings is that the sub-space with optimal values in the soft-prompt space has probably a convex shape. This may be indicated by the fact that task prompts trained from different random initializations for the same task do not point to the same direction (based on Figures 2 and 3), but still achieve identical results.
Section 4.3 implied that we can use different combinations of task prompt vectors to gain even zero-shot multi-task behavior (in Figure 4). We can combine multiple task prompt vectors and maintain multi-task performance on the source tasks, but the right task combinations need to be found (e.g., by evaluating on held-out validation sets). We can see that it is possible to use linear combinations even though the soft-prompts space is non-linear. The rationale behind this could be that task prompt vectors are linear approximations of how the soft-prompts change during the training. Another possibility may be that the task prompt vectors are sparse, and a combination of 2 sparse task prompt vectors creates a vector that contains more information about both tasks. These findings can be further useful for machine unlearning tasks, where one could also include subtraction.
@@ 6 Conclusion
In our work, we introduce and investigate task prompt vectors as a method of multi-task transfer from prompt tuning. We show that the task prompt vectors are not dependent on random initialization and that the performance across different random initializations does not change significantly in the majority of observed source tasks.We show that in some tasks, the combination via arithmetic addition maintains the single-task performance. Finally, we show that certain combinations of task prompt vectors can be a better option for initialization for certain tasks while maintaining higher multi-task modularity than other methods.
In the future, we would like to evaluate the cross-model performance of task prompt vectors. We think that further experiments with generation tasks may be another interesting extension. Moreover, task prompt vector arithmetic has the highest potential for improving the unlearning in PLMs by negating the task prompt vectors for the tasks we want to unlearn. Such an option is enabled by introducing task prompt vectors, which would not be possible with the existing state-of-the-art methods.
@@ Acknowledgements
The experiments in our work utilize only classification datasets in the English language with various characteristics (problem type, dataset size, number of classes), mostly because of the large amount of publicly available data covering a variety of NLU problems. At the same time, to direct our focus primarily on the evaluation of task prompt vectors, we utilize only monolingual models in the scope of our work.
Even though there are many other PLMs capable of conditional generation that beat T5 models in performance on various benchmarks, we focus our experiments on the T5-base model as it is commonly used as a representative model in many PEFT methods. We also utilize LLaMa-3.1-8B-Instruct model only for the initialization dependency experiments.
Finally, we focus on 3 common NLU problems (natural language inference, topic classification, and sentiment classification) that are commonly incorporated in NLU benchmarks and do not consider other NLU problems (e.g., question answering, slot tagging, acceptability classification, text generation). However, we find that our set of 3 common NLU problems each covering 4 different tasks, is enough to evaluate the properties of task prompt vectors. Including generation tasks would further introduce a problem with the evaluation of the generated text.
@@
The experiments in this paper were conducted with publicly available datasets MNLI, QNLI, SciTail, SNLI, DBPedia, TREC, AG News, Yahoo Answers, SST2, Yelp Polarity, SST5, and IMDB, citing the original authors. MNLI, QNLI, and SST2 are part of the GLUE benchmark. As we were not able to determine the license for all used datasets, we have opted to use them as in a limited form as possible, adhering to the terms of use of the GLUE benchmark for all of the mentioned datasets. As the datasets are commonly used in other related works, and were published in scientific works that went through an established review process, we do not check for the presence of any offensive content as it was already removed by the authors of these publicly available datasets. In addition, we do not utilize any personally identifiable information or offensive content and we do not perform crowdsourcing in any form for data annotation. To our knowledge, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the field of Machine Learning (i.e., use of computational resources that are consuming energy and producing heat with indirect CO2 emission production). We follow the license terms for the T5-base model we use – all models and datasets allow their use as part of the research. As we perform conditional generation transform into the classification problem (generating only labels), we minimize the problem of generating offensive or biased content.
@@
The experiments in this paper require a significant amount of GPU computing resources as we train and evaluate 1 model over multiple random initializations (10) for different methods (4) and datasets (12). Overall the experiments including evaluations (which did not require training, but still used GPU resources for inference) and preliminary experiments (which are not reported in the scope of our work) were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh. Approximately, 1000 hours of computation performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 105.24 kgCO2eq of which 0 percent were directly offset. These estimations were conducted using the CodeCarbon Courty et al. (2024) python module. Whenever possible we tried to reduce the computational costs. Because our method is built upon the prompt tuning PEFT method, we always trained only a small part of the model parameters (76800 parameters, which is around 0.2% of the T5-base model parameters), and training the model fully will probably require more GPU hours and create more CO2 emissions.
@@ Appendix A Experimental setup: Further Details
We provide all of our configurations in the config directory of our repository. We set soft-prompt length to 100 tokens, learning rate to 0.3, and lower the weight decay of the AdamW optimizer (Loshchilov and Hutter, 2019) to $1\times 10^{-5}$ for T5-base model. We also utilize the Seq2SeqTrainer class from the Huggingface transformers Python module. We train all models on all data sets for 10 epochs, except for TREC, where we train for 50 epochs due to the tendency of models to underfit here. We set different hyperparameters for prompt tuning and the hyperparameters for zero- or few-shot evaluation do not differ much from the hyperparameters for prompt tuning. We train for 1000 update steps while keeping a batch size of 2 for 5, 10, 25 shots, 8 for 50, 100, 250 shots, and 16 for 500, 750, 1000 shots. In general, we chose the maximum token length for labels by searching the dataset for the maximum token length (in our configs, we set default max_target_lenght to 128 if the dataset requires to generate sentences), for the inputs we pad the token sequences to 256 tokens with the max_target_lenght parameter. We use a learning rate of 0.3 for the AdamW optimizer, with weight decay of $1\times 10^{-5}$ and 500 warmup steps for 10 epochs (with an exception for the TREC dataset) with the batch size of 32. We evaluate, log and save after each 100 training steps and keep only the best model at the end of the training. In our configs, we set a number of tokens to 50, but in reality, Hugging Face peft library doubles the number for encoder-decoder models like T5. When combining task prompt vectors, we evaluate their performance on the individual source tasks that formed the task combination and found the best rescaling factor $\lambda$ via held-out validation sets.
For training the soft-prompts with the LLaMa-3.1-8B-Instruct model we utilized similar hyperparameter settings as for the T5-base model with the exception of using the cosine function for the learning rate scheduler and the usage of the SFTTrainer class from the Huggingface trl Python module von Werra et al. (2020) for training.
Since we are utilizing the T5-base for conditional generation, we are computing exact match instead of accuracy for classification. Because we are generating labels also for classification tasks, the exact match is equivalent to accuracy in the sequence classification task. In the scope of our work, we refer to a single dataset as a task.
@@ Appendix B Additional results: Task Prompt Vectors and Task Prompt Cosine Similarities
Figure 6: Comparisons of cosine similarities of task prompts fine-tuned on different tasks for T5-base model. Each heatmap represents a different task combination. We calculate the cosine similarities for all combinations of 10 random initializations omitting the combinations of random initializations where cosine similarity is equal to 1 (single-task comparisons). Each heatmap is represented as a single field in Figure 2 by averaging all values. The x and y axes represent the number of random initializations.
@@
Figure 8: Comparisons of average cosine similarities of task prompt vectors for T5-base model. The averages are calculated similarly to Figure 6 but with task prompt vectors created from different task prompts. Each heatmap represents a different task combination. We calculate the cosine similarities for all combinations of 10 random initializations omitting the combinations of random initializations where cosine similarity is equal to 1 (single-task comparisons). Each heatmap is represented as a single field in Figure 3 by averaging all values. The x and y axes represent the number of random initializations.
@@
Figure 11: Comparison of average cosine similarities of task prompts fine-tuned on different tasks for LLaMa-3.1-8B-Instruct model. The average is calculated across all combinations of 3 random initializations (i.e., row QNLI column MNLI was calculated as the average of all cosine similarities between MNLI and QNLI task prompts for all random initialization combinations omitting the combinations where cosine similarity is equal to 1). The diagonal represents the cosine similarities of the same tasks and it represents the maximum value of cosine similarity across different random initializations.
Figure 12: Comparison of average cosine similarities of task prompt vectors. The averages are calculated equivalently to Figure 11 but with task prompt vectors created from different task prompts.